OpenStudy (anonymous):

Hey guys if you can help me I will give you a medal and fan you! I need help with linear application!!!

OpenStudy (anonymous):

8x+2y=4 -4x+3y=-10 Solve with linear application (with or without multiplication)

OpenStudy (anonymous):

A. (1,3) B. (2,3) C. (1,2) D. (1,-2)

OpenStudy (anonymous):

u can plug in each answer nd try to c if they make it true

OpenStudy (anonymous):

But how do I do that?

OpenStudy (anonymous):

So do you think it would be d.? @rocker128

OpenStudy (anonymous):

YES :)

OpenStudy (anonymous):

Can you help me with another one?

OpenStudy (anonymous):

I have 3 questions I dont understand at all!

OpenStudy (anonymous):

SURE ill do my best

OpenStudy (anonymous):

Ok great thanks so much!

OpenStudy (anonymous):

4g+5h=4 8g-10h=4

OpenStudy (anonymous):

what do they want u to do here?

OpenStudy (anonymous):

A. (3/4, -1/5) B. (-3/4, -1/5) C. (3/4,1/5) D. (-3/4, 1/5)

OpenStudy (anonymous):

i have to solve using linear combination with or without mulitplication

OpenStudy (anonymous):

ahh same thing. just plug in

OpenStudy (anonymous):

How do I do that with fractions?

OpenStudy (anonymous):

u do the same thing with fractions nothing changes

OpenStudy (anonymous):

ok so let me work it out and then can you check it?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

So is it c.?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

Ok can I do the other 2 write them down and you can check them?

OpenStudy (anonymous):

PRETTY PLEASE????

OpenStudy (anonymous):

surrre

OpenStudy (anonymous):

x+y=50 0.05x+0.06y=2.8 A.(20,-50) B.(10,30) C. (40,60) D. (20,30)

OpenStudy (anonymous):

is it D.?

OpenStudy (anonymous):

@rocker128

OpenStudy (anonymous):

5x-6y=27 3x+2y=-23 A. (9,4) B. (-3,-7) C. (3,7) D. (3,-7) This is the next question but was the one above D.?

OpenStudy (anonymous):

the one before is d